AS/AD Model.
AD: P = 90 - 3Y.
AS: P = 15.
PO: Y = 30.
What is the Inflationary Gap?
1
Expert's answer
2016-05-14T12:37:03-0400
Find macroeconomic equilibrium, when total demand=total supply
P=90-3Y=15
Y (actual GDP) =25
When actual GDP is below the potential GDP the economy has negative output gap, also called as recessionary gap.
Inflationary Gap = Actual GDP- Potential GDP=25-30=-5
We received negative inflationary gap, also called as recessionary gap.
The inflationary gap equals minus 5.